OpenStudy (anonymous):

Write an equation for the line in point-slope and slope-intercept forms. 4) Passes through (1, 2) and has a slope of 7. Plz help can someone plz show me how to do this??????!!!!!!

OpenStudy (zale101):

\[m=\frac{ y-y _{1} }{ x-x _{1} }\]

OpenStudy (zale101):

use this formula

OpenStudy (zale101):

in which m is the slope and the values of \[x_{1}\] and \[y_{1}\] come from the x and y values of the point you were given.

OpenStudy (zale101):

Can u Substitute those values

OpenStudy (anonymous):

wait substitute them how?

OpenStudy (zale101):

plug in to the equations

OpenStudy (anonymous):

but I only have one x and y

OpenStudy (zale101):

m is 7 x is 1 y is 2

OpenStudy (anonymous):

ok so can my other point be (2,18)?

OpenStudy (zale101):

now plug in these numbers to the formula i showed u above

OpenStudy (zale101):

Another way you can do it is to use the slope-intercept form of an equation.

OpenStudy (zale101):

y=mx+b

OpenStudy (anonymous):

ok so what's my answer?

OpenStudy (zale101):

y=mx+b where m is the slope of the line b is the value where the graph crosses the y axis. we already know that 7 is our slope so y=7x+b

OpenStudy (zale101):

Substitute x = -1 and y = -2 values into the slope intercept form.

OpenStudy (zale101):

can u do this?

OpenStudy (zale101):

i will do it for u 2=7*1+b

OpenStudy (zale101):

understood?

OpenStudy (anonymous):

so my equation for the line in point slope, and slope intercept forms that passes through (1,2) and has a slope of 7 is 2=7*1+b?

OpenStudy (zale101):

correct